OpenStudy (anonymous):

How many solutions does the equation have? 3(4x-3)-7x=5x-9 A; One solution B: infinite solutions C:No solution

OpenStudy (anonymous):

@rational

OpenStudy (er.mohd.amir):

since eq is linear so solution is one.that is x power is one .

OpenStudy (anonymous):

can u help me with more?

OpenStudy (er.mohd.amir):

yes on what

OpenStudy (anonymous):

Which expressions would complete this equation so that it has one solution? 4(2x – 3) – 4x = Choose exactly two answers that are correct. A;4x – 12 B:3x – 4 c: 3(2x – 6) d:4(x – 5)

OpenStudy (er.mohd.amir):

first solve this and find value of x

OpenStudy (er.mohd.amir):

actually first Q is an identity so its holds good for all value of x so ans is infinite solution

OpenStudy (er.mohd.amir):

for second one option A gives an identity 4(2x-3)-4x=8x-12-4x=4x-12 so D also vanish x

OpenStudy (anonymous):

so the answer is a and d?

OpenStudy (er.mohd.amir):

no ans is B and C ,since A and D give infinite solution